OpenStudy (anonymous):

dy/dx= 3 cos 2x -(2/x^2)

OpenStudy (anonymous):

\[dy = 3 \cos 2x dx - 2 \times x^{-2} dx\]

OpenStudy (anonymous):

Now integrating both sides \[\int dy = 3 \int \cos 2x dx - \int 2 \times x^{-2} dx\]